# arXiv:2602.14013v2[math.CV] 01 Mar 2026

### v2

@@ 6 The versions for prime ends
Assume the contrary. Then there exists $a>0$ such that for $\delta=1/k,$ $k=1,2,\ldots,$ there are $m_{k}\in{\mathbb{N}}$ and $x_{k},x^{\,\prime}_{k}\in D_{m_{k}}$ such that $\rho(x_{k},x^{\,\prime}_{k})<1/k,$ however, $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a.$ By the definition, $\overline{D_{0}}_{P}$ is a compact space, so that we may consider that the sequences $x_{k},x^{\,\prime}_{k}\rightarrow P_{0}\in\overline{D_{0}}_{P}$ by the metrics $\rho.$ Arguing similarly to the proof of Lemma 6, we may consider that the sequence $m_{k}$ is increasing by $k,$ besides that, we may consider that $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a$ holds for $m$ instead of $m_{k},$ i.e., $x_{m},x^{\,\prime}_{m}\in D_{m}$ and (6.8) holds. Now, $f_{m}$ is equicontinuous at $x_{0}$ (see [RS, Lemma 4.9] for $p=n$ and [Sev2, Lemma 4.3] for n1<p<n.n-1<p<n. Now, we may consider that $P_{0}\in E_{D_{0}}=\overline{D_{0}}_{P}\setminus D_{0}.$
Let $\sigma_{m},$ $m=0,1,2,\ldots,$ lie on the spheres $S(z_{0},r_{m})$ centered at some point $z_{0}\in\partial D_{0},$ where $r_{m}\rightarrow 0$ as $m\rightarrow\infty$ (such a sequence $\sigma_{m}$ exists by [IS, Lemma 3.1], cf. [KR]). Since $D_{m}$ is a regular sequence of domains with a respect to $D_{0}$ in the terms of prime ends, there exists a sequence of cuts $\varsigma_{m}$ and domains $d^{\,\prime}_{m},$ equivalent to $\sigma_{m}$ and $d_{m},$ respectively, such that the following condition holds: given $d^{\,\prime}_{k},$ $k=1,2,\ldots,$ there is $M_{1}=M_{1}(k)$ such that $d^{\,\prime}_{k}\cap D_{m}$ is a non-empty connected set for every $m\geqslant M_{1}(k).$ Now, we will consider that $d^{\prime}_{k}\subset d_{k}$ for any $k=1,2,\ldots.$
Arguing similarly to the proof of Lemma 3, we may consider that $x_{m},x^{\,\prime}_{m}\in d^{\prime}_{m}\cap D_{m}$ while $d^{\prime}_{m}\cap D_{m}$ is connected. We join the points $x_{m}$ and $x^{\,\prime}_{m}$ by a path $\gamma_{m}:[0,1]\rightarrow D_{m}\cap d^{\prime}_{m}$ such that $\gamma_{m}(0)=x_{m},$ $\gamma_{m}(1)=x^{\,\prime}_{m}$ and $\gamma_{m}(t)\in D_{m}\cap d^{\prime}_{m}$ for $t\in(0,1).$
@@
By the definition of the family $\mathfrak{R}_{Q,\delta,p,E}(D_{0},\mathfrak{D}),$ for any $f_{m}\in\mathfrak{R}_{Q,\delta,p,E}(D_{0},\mathfrak{D})$ and any domain $D^{\,\prime}_{m}:=f_{m}(D_{m})$ there exists a continuum $K_{m}\subset D^{\,\prime}_{m}$ such that $h(K_{m})\geqslant\delta$ and $h(f_{m}^{\,-1}(K_{m}),\partial D_{m})\geqslant\delta>0.$ Since by the condition of the lemma the domains $D^{\,\prime}_{m}$ form the equi-uniform family with respect to $p$-modulus and due to (2.6), we obtain that the inequality
$M_{p}(\Gamma(K_{m},C_{m},D^{\,\prime}_{m}))\geqslant b$ (6.13)
@@
holds for all $m=1,2,\ldots$ and some $b>0.$ Consider the family $\Gamma_{m},$ consisting of all paths $\beta:[0,1)\rightarrow D^{\,\prime}_{m},$ where $\beta(0)\in C_{m}$ and $\beta(t)\rightarrow p\in K_{m}$ for $t\rightarrow 1.$ Let $\Gamma^{*}_{m}$ be a family of all total $f_{m}$-liftings $\alpha:[0,1)\rightarrow D$ of the family $\Gamma_{m}$ with origin on $\gamma_{m}$ (see Proposition 3). Since $f_{m}$ is closed, we have: $\alpha(t)\rightarrow f^{\,-1}_{m}(K_{m}),$ where $f^{\,-1}_{m}(K_{m})$ is the total preimage of the continuum $K_{m}$ under the mapping $f_{m}.$ Since $h(f_{m}^{\,-1}(K_{m}),\partial D_{m})\geqslant\delta>0$ and $D_{m}\subset D_{0},$ by Proposition 6, $h(f_{m}^{\,-1}(K_{m}),\partial D_{0})\geqslant\delta>0,$ as well. Now, we may consider that $f_{m}^{\,-1}(K_{m})\subset D_{0}\setminus d^{\,\prime}_{1}.$ In this case, by Proposition 3
@@
for each measurable function $\eta:(r_{1},r_{m})\rightarrow[0,\infty],$ such that $\int\limits_{r_{1}}^{r_{m}}\eta(r)\,dr\geqslant 1.$ Note that, the function
[…]
where $I(r_{m},r_{1}):=\int\limits_{r_{m}}^{r_{1}}\psi(t)\,dt,$ satisfies the normalization condition (1.4). Therefore from the conditions (2.3)–(2.4) and (6.15) it follows that
@@
Funding. The paper was supported by the National Research Foundation of Ukraine, Project number 2025.07/0014, Project name: ‘‘Modern problems of Mathematical Analysis and Geometric Function Theory’’.

### v3

@@ 6 The versions for prime ends
Assume the contrary. Then there exists $a>0$ such that for $\delta=1/k,$ $k=1,2,\ldots,$ there are $m_{k}\in{\mathbb{N}}$ and $x_{k},x^{\,\prime}_{k}\in D_{m_{k}}$ such that $\rho(x_{k},x^{\,\prime}_{k})<1/k,$ however, $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a.$ By the definition, $\overline{D_{0}}_{P}$ is a compact space, so that we may consider that the sequences $x_{k},x^{\,\prime}_{k}\rightarrow P_{0}\in\overline{D_{0}}_{P}$ by the metrics $\rho.$ Arguing similarly to the proof of Lemma 6, we may consider that (6.1) holds. If $P_{0}\in D_{0}$, then $f_{m}$ is equicontinuous at $P_{0}$ (see [RS1, Lemma 4.9] for $p=n$ and [Sev1, Lemma 4.3] for n1<p<n.n-1<p<n. Now, we may consider that $P_{0}\in E_{D_{0}}=\overline{D_{0}}_{P}\setminus D_{0}.$ Under notations of Lemma 6 we may consider that $x_{m},x^{\,\prime}_{m}\in d^{\,\prime}_{m}\cap D_{m}$ while $d^{\,\prime}_{m}\cap D_{m}$ is connected. We join the points $x_{m}$ and $x^{\,\prime}_{m}$ by a path $\gamma_{m}:[0,1]\rightarrow D_{m}\cap d^{\,\prime}_{m}$ such that $\gamma_{m}(0)=x_{m},$ $\gamma_{m}(1)=x^{\,\prime}_{m}$ and $\gamma_{m}(t)\in D_{m}\cap d^{\,\prime}_{m}$ for $t\in(0,1).$ Set $C_{m}:=|f_{m}(\gamma_{m})|.$ Now, $h(C_{m})\geqslant a/2$ for sufficiently large $m\in{\mathbb{N}}.$
@@
By the definition of $\mathfrak{R}_{Q,\delta,p,E}(D_{0},\mathfrak{D}),$ for any $f_{m}\in\mathfrak{R}_{Q,\delta,p,E}(D_{0},\mathfrak{D})$ and any domain $D^{\,\prime}_{m}:=f_{m}(D_{m})$ there exists a continuum $K_{m}\subset D^{\,\prime}_{m}$ such that $h(K_{m})\geqslant\delta$ and $h(f_{m}^{\,-1}(K_{m}),\partial D_{m})\geqslant\delta>0.$ Since by the assumption the domains $D^{\,\prime}_{m}$ form the equi-uniform family with respect to $p$-modulus and $h(C_{m})\geqslant a/2$ we obtain that
$M_{p}(\Gamma(K_{m},C_{m},D^{\,\prime}_{m}))\geqslant b\,$, (6.6)
@@
$m=1,2,\ldots,$ for some $b>0.$ Consider the family $\Gamma_{m},$ consisting of all paths $\beta:[0,1)\rightarrow D^{\,\prime}_{m},$ where $\beta(0)\in C_{m}$ and $\beta(t)\rightarrow p\in K_{m}$ as $t\rightarrow 1.$ Let $\Gamma^{*}_{m}$ be a family of all total $f_{m}$-liftings $\alpha:[0,1)\rightarrow D$ of the family $\Gamma_{m}$ with origin at $|\gamma_{m}|$ (see Proposition 3). Since $f_{m}$ is closed, $\alpha(t)\rightarrow f^{\,-1}_{m}(K_{m}),$ where $f^{\,-1}_{m}(K_{m})$ is the total preimage of the continuum $K_{m}$ under the mapping $f_{m}.$ Since $h(f_{m}^{\,-1}(K_{m}),\partial D_{m})\geqslant\delta>0$ and $D_{m}\subset D_{0},$ by Proposition 4, $h(f_{m}^{\,-1}(K_{m}),\partial D_{0})\geqslant\delta>0,$ as well. Now, we may consider that $f_{m}^{\,-1}(K_{m})\subset D_{0}\setminus d^{\,\prime}_{1}.$ In this case, by Proposition 3
@@
for each measurable function $\eta:(r_{m},r_{1})\rightarrow[0,\infty],$ such that $\int\limits_{r_{1}}^{r_{m}}\eta(r)\,dr\geqslant 1.$ Again, arguing similarly to the proof of Lemma 6, we obtain from (6.8) that
@@
as $m\rightarrow\infty,$ where $\alpha(\varepsilon)$ is some nonnegative function tending to zero as $\varepsilon\rightarrow 0$ (which exists by the condition (2.2)). In addition, $f_{m}(\Gamma^{\,*}_{m})=\Gamma_{m}$ and $M_{p}(\Gamma_{m})=M_{p}(\Gamma(K_{m},C_{m},D^{\,\prime}_{m})).$ Thus
@@
Example 1. Let $D_{0}$ be the unit disk $\mathbb{D}$} in $\mathbb{C}$} with a cut along the segment $[0,1]$, i.e., $D_{0}={\mathbb{D}}\setminus I\subset{\mathbb{C}}$, where ${\mathbb{D}}:=\{z\in{\mathbb{C}}\,:\,|z|<1\}$, $I:=\{z=x+iy\in{\mathbb{C}}\,:\,y=0,0\leqslant x\leqslant 1\}$. Let $f_{m}$ be a conformal mapping of $D_{m}=\{z\in{\mathbb{C}}\,:\,|z|<m/(m+1)\}\setminus I_{m},$ $I_{m}:=\{z=x+iy\in{\mathbb{C}}\,:\,y=0,0\leqslant x\leqslant m/(m+1)\},$ onto $\mathbb{D}$}. For instance, we may take
$f_{m}(z)=\frac{\left(\frac{\sqrt{zr_{m}}+1}{\sqrt{zr_{m}}-1}\right)^{2}-i}{\left(\frac{\sqrt{zr_{m}}+1}{\sqrt{zr_{m}}-1}\right)^{2}-i}\,,\quad z=x+iy\,,\quad i^{2}=-1\,$,
where $\sqrt{z}$ denotes the main branch of the square root and $r_{m}=(m+1)/m.$ We verify conditions of Theorem 1. First of all, $f_{m}$ are homeomorphisms of $D_{m}$ onto $\mathbb{D}$} satisfying (1.3)–(1.4) at any point of $\overline{D_{0}}=\overline{\mathbb{D}}$ with $Q\equiv 1$ and $p=n=2$ as conformal mappings (see [Va, Theorem 8.1]). Since this sequence (obviously) converges to the conformal mapping $f(z)=\frac{\left(\frac{\sqrt{z}+1}{\sqrt{z}-1}\right)^{2}-i}{\left(\frac{\sqrt{z}+1}{\sqrt{z}-1}\right)^{2}-i}$ locally uniformly in $D_{0},$ the relation $h(f_{m}(A))\geqslant\delta$ obviously holds for any $m=1,2,\ldots,$ any continuum $A$ in $D_{0}\cap\bigcap\limits_{m=1}^{\infty}D_{m}$ and some $\delta>0.$ Besides that, the relation $h(\overline{{\mathbb{R}}^{n}}\setminus f_{m}(D_{m}))=h(\overline{{\mathbb{R}}^{n}}\setminus{\mathbb{D}})\geqslant\delta$ holds, for in0stance, with $\delta=1.$ Observe that $D_{m},$ $m=1,2,\ldots,$ converge to $D_{0}={\mathbb{D}}\setminus I$ as its kernel, see Figure 4.
Figure 4: Illustration for Example 6.
Besides that, the domains $d_{k},$ formed by chains of cuts lying on the circles centered at some point $z_{0}\in\partial{\mathbb{D}}\cup I$ form a connected intersection with the domains $D_{m}$ for each fixed $k$ and all sufficiently large $m.$ Indeed, prime ends of $D_{0}$ may be conditionally divided into 6 groups: 0) the set of prime ends $P_{0}$ which correspond to $z_{0}\in{\mathbb{S}}^{1}\setminus I,$ where ${\mathbb{S}}^{1}=\partial{\mathbb{D}};$ 1) and 2): the prime ends $P_{1}$ and $P_{2}$ with impression $z_{0}=1,$ the sequences of domains $d_{k}$ lie from below and from upper of $I,$ correspondingly; 3) and 4): the set of prime ends $P_{3}$ and $P_{4}$ with impression at some $z_{0}\in I\setminus\{1,0\},$ the sequences of domains $d_{k}$ lie from below and from upper of $I,$ correspondingly; 5) the prime end $P_{5}$ which correspond to its impression $I(P_{5})=\{0\}.$ Let $d_{k},$ $k=1,2,\ldots,$ be a sequence of domains in $P_{i},$ $i=\overline{0,5}.$ Now, $d_{k}$ is convex for any $k=1,2,\ldots.$ Given $k\in{\mathbb{N}},$ the intersection $D_{m}\cap d_{k}$ is non-empty and convex for sufficiently large $m.$ In particular, $D_{m}\cap d_{k}$ is connected for above $m,$ as required. Thus, 1) the sequence of domains $D_{m}$ is regular with respect to $D_{0}$ in terms of prime ends. Observe that, 2) for every $m\in{\mathbb{N}},$ a domain $D_{m}$ is quasiconformally regular; $D_{0}$ is also quasiconformally regular. Indeed, $D_{0}$ and $D_{m}$ are regular because by the Riemannian mapping theorem, $D_{0}$ (or $D_{m}$) is conformally equivalent to $\mathbb{D}$}; moreover, $\mathbb{D}$} has locally quasiconformal boundary (see e.g. [Va, Theorem 17.10]). Note that, 3) the family $f_{m}(D_{m})={\mathbb{D}}$ is equi-uniform with respect to $2$-modulus over all $m\in{\mathbb{N}}$ (see Remark 1). Finally, 4) $Q\equiv 1$ has a finite mean oscillation in $\overline{D_{0}};$ moreover, (1.5) holds for every $x_{0}\in\overline{D_{0}}$ and any $\beta(x_{0})>0.$ Thus, all of conditions of Theorem 1 hold, as required.
Example 2. It is possible to construct corresponding families of mappings, satisfying Theorem 1 that have unbounded characteristics. Let $z_{0}\in D_{0}$ and $0<r_{0}<d(z_{0},\partial D_{0}),$ where $D_{0}$ is a domain from Example 6. Put $h(x)=\frac{x}{|x|\log\frac{(r_{0}e)}{|x|}},$ $x\in B(z_{0},r_{0}),$ $h(z_{0})=z_{0},$ $h|_{S(z_{0},r_{0})}=x.$ Then $h$ is defined in the ball $B(z_{0},r_{0})$ and $h(B(z_{0},r_{0}))=B(z_{0},r_{0}).$ Reasoning similarly to [MRSY, Proposition 6.3], it may be shown that $h$ satisfies the relations (1.3)–(1.4) at any point $x_{0}\in\overline{B(z_{0},r_{0})}$ for $Q=Q(x)=\log\left(\frac{r_{0}e}{|x|}\right).$ Note that $Q$ satisfies (1.5) for every $x_{0}\in\overline{D_{0}}$ and any $\beta(x_{0})>0$ that may be verified directly. Now, we put
$g_{m}(x)=\begin{cases}(f_{m}\circ h)(x)\,,&x\in B(z_{0},r_{0})\,,\\ f_{m}(x)\,,&x\not\in B(z_{0},r_{0})\end{cases}\,$,
where $f_{m}$ and $f$ are from Example 6. By the construction, $g_{m}$ satisfy the relations (1.3)–(1.4) at any point $x_{0}\in\overline{D_{0}}$ for $Q=Q(y)=\log\left(\frac{r_{0}e}{|x|}\right).$ Observe that, the domains $f_{m}(D_{m})$ are not changed, so that the domains $g_{m}(D_{m})$ are regular, as well. In addition, since $h$ is a fixed mapping, the relations $h(g_{m}(A))\geqslant\delta,$ $m=1,2,\ldots,$ also hold for some $\delta>0$ and all $m=1,2,\ldots.$ The mappings $g_{m},$ $m=1,2,\ldots,$ are homeomorphisms and satisfy all the conditions of Theorem 1.
Declarations.
Conflict of interest. The authors have no financial or proprietary interests in any material discussed in this article.
Funding. The work was supported by the National Research Foundation of Ukraine, Project number 2025.07/0014, Project name: ‘‘Modern problems of Mathematical Analysis and Geometric Function Theory’’.
Data availability statement. All necessary data are included in the paper.
Contribution. The authors have equal contribution to writing the article.